OpenStudy (anonymous):

determine whether the graphs of each pair of lines are parallel 3x+4=7 2y=6x-7

OpenStudy (anonymous):

your missing a variable in the 1st equation

OpenStudy (anonymous):

Assuming 4y for the second term in the first equation

OpenStudy (anonymous):

They are not parallel since the slope are different

OpenStudy (anonymous):

\[m _{1}=-3/4\]and\[m _{2}=3\]

OpenStudy (anonymous):

Write both equations in the form of y = mx + b. That means to solve for y in terms of everything else. Then compare the m of the two equations if they are the same then the two lines are parallel if they are negative inverse of each other the lines or orthogonal and if they are completely different (unrelated) the lines do not intersect.

OpenStudy (anonymous):

im sorry it is 3x+4=y 2y=6x-7

OpenStudy (anonymous):

So for the first equation it's already in solve intercept form: y = 3x + 4...(1) For the second equation we have to divide both sides by 2 so that we are left with only a y on the left hand side: y = 3x - 7/2...(2)

OpenStudy (anonymous):

So you compare equations (1) and (2) and you realize that indeed the slopes with are 3 are the same therefore the two lines are parallel.

OpenStudy (anonymous):

If you want to do a comparison make sure to write the equation of the line in slope-intercept form: y = mx + b. Then you can easily identify the slope of the line which is m(coefficient of x). Ideally you also want to get in the habit of representing the equation of lines in this standard form.